A job with a 95% learning curve required 19 hours for the second unit. The first unit should require approximately how many hours?
Answer #1

Tn = T1*n^(ln(l)/ln(2)) Where Tn = time required to do the nth Job

Here t2 = 19 hours and l = learning rate = 95% = 0.95, n = 2

hence, 19 = T1*2^(ln(0.95)/ln(2))

or, T1 = 19/(2^(ln(0.95)/ln(2))) = 20 hours = Time required for the first unit (Ans)